# Boundary Attention: Learning to Localize Boundaries under High Noise

### v2

@@ 7 Conclusion
Fig. S6: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [19] and Pidinet [23] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [3], and Field of Junctions [25] with patch size $11$.
@@
In Figures S7 and S6, we show how the model behaves on noiseless natural images that contain texture and recognizable objects. In particular, Figure S6 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations. The figure compares our output to that from Canny [3], Field of Junctions [25] with a patch size of $11$, Pidinet [23], and EDTER [19]. The latter two methods are trained on human annotated data, whereas the former two methods, like our model, are not. (Note that inputs for all models besides EDTER [19] were $300\times 400$. Input to EDTER was down-sampled to $225\times 300$ due to its input size constraint.)
@@
We find that our model produces finer structures than the end-to-end learned models [7, 19] because it is trained to only use local spatial averages of color as its cue for boundaries and grouping. It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans. Compared to the bottom-up methods of Canny [3] and Field of Junctions [25], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of windowing field $\mathbf{p}[k]$. In contrast, the Field of Junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
@@
Figure S9 provides additional comparisons for a sample of varying-noise images from the ELD dataset [26]. When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases. Some algorithms, such as Field of Junctions [25], have tunable parameters such as patch-size that provide control over the level of detection. A small patchsize allows recovering fine structures in lower noise situations, but it causes many false positive boundaries at high noise levels. Conversely, a large patchsize provides more resilience to noise but has not ability to recover fine structure at all. Our model reduces the severity of this trade-off by automatically adapting its local windowing functions in ways that have learned to account for both the amount of noise and the local geometry of the underlying boundaries.
In Figure S9 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods. We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—UAED [28], EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve. We emphasize again that a user can adjust the behavior of Canny and Field of Junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for Field of Junctions. Increasing the local size improves their resilience to noise but reduces their spatial precision. Neither system provides the ability to estimate fine grained details and withstand noise, like our model does.
@@
Fig. S9: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise. We compare to end-to-end models that are trained to match human annotations (UAED [28], EDTER [19], HED [27], Pidinet [23], and Structured Edges (SE) [7]) in addition to low-level models that are not (Canny [3], gPb [14], and Field of Junctions (FoJ-17) [25]) with patch size 17.

### v3

@@ 7 Conclusion
Fig. S3: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [18] and Pidinet [21] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [2], and the field of junctions (FoJ) [23] with patch size $11$.
@@
In Figures S4 and S3, we show how the model behaves on noiseless natural images that contain texture and recognizable objects. In particular, Figure S3 emphasizes how the boundary maps produced by our model qualitatively differ from other methods. Here, in addition to showing the results those of classical bottom-up edge-detectors, we include results learned, end-to-end models that have been trained to match human annotations as another point of reference. It is important to remember that these methods are trained to identify boundary structures that are defined by semantic variations, whereas our method and other low-level methods divide regions based on variations in color.
Figure  S3 compares our output to that from Canny [2], the field of junctions (FoJ) [23] with a patch size of $11$, Pidinet [21], and EDTER [18], the latter two being networks trained on human annotated data. (Note that inputs for all models besides EDTER [18] were $300\times 400$. Input to EDTER was down-sampled to $225\times 300$ due to its input size constraint.)
@@
We find that our model produces finer structures than the end-to-end learned models [6, 18] because it is trained to only use local spatial averages of color as its cue for boundaries and grouping. It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans. Compared to the bottom-up methods of Canny [2] and the field of junctions [23], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of windowing field $\mathbf{p}[k]$. In contrast, the the field of junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
@@
Figure S6 provides additional comparisons for a sample of varying-noise images from the ELD dataset [24]. We include results from other low level methods and as a point of comparison, the outputs of several methods trained to parse semantic boundaries.
When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases. Some algorithms, such as the field of junctions (FoJ) [23], have tunable parameters such as patch-size that provide control over the level of detection. A small patchsize allows recovering fine structures in lower noise situations, but it causes many false positive boundaries at high noise levels. Conversely, a large patchsize provides more resilience to noise but has not ability to recover fine structure at all. Our model reduces the severity of this trade-off by automatically adapting its local windowing functions in ways that have learned to account for both the amount of noise and the local geometry of the underlying boundaries.
In Figure S6 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods. We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—UAED [27], EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve. We emphasize again that a user can adjust the behavior of Canny and the field of junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for the field of junctions. Increasing the local size improves their resilience to noise but reduces their spatial precision. Neither system provides the ability to estimate fine grained details and withstand noise, like our model does.
@@
Fig. S6: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise. We compare to end-to-end models that are trained to match human annotations (UAED [27], EDTER [18], HED [25], Pidinet [21], and Structured Edges (SE) [6]) in addition to low-level models that are not (Canny [2], gPb [13], and the field of junctions (FoJ-17) [23]) with patch size 17.